# FCMR: Robust Evaluation of Financial Cross-Modal Multi-Hop Reasoning

### v1

@@ 1 Introduction
Despite the recent progress in AI (Touvron et al., 2023; Anthropic, 2024; OpenAI, 2024; inter alia), developing systems capable of human-level reasoning remains a significant challenge. Human cognition involves integrating information from multiple modalities to comprehend and make decisions in the real world. A domain that primarily requires such a comprehensive understanding is finance, where analysts often need to simultaneously examine textual reports, tabular data (e.g., balance sheets), and visual data (e.g., charts). For example, verifying the statement from Figure 2“The corporation, with the smallest act value in the years when the fopo value for ABBOTT LABORATORIES is below 730.5, is entitled to receive \$43 million in minimum sublease income from non-cancelable subleases.”—one must consider all the relevant clues provided by each source, an ability we refer to as cross-modal multi-hop reasoning.
@@
Second, current benchmarks are largely focused on testing straightforward problems, such as single- and two-hop reasoning. As shown in Figure 1, MMQA Talmor et al. (2021)—one of the leading benchmarks in this field—features only about 0.8% of the queries that explicitly require three-hop cross-modal reasoning. Furthermore, in preliminary experiments, we discovered that GPT-4o OpenAI (2024) can solve the MMQA’s most challenging problems with a 43% exact match accuracy, even without access to visual clues. This result highlights the urgent need to establish a higher standard for evaluating cross-modal multi-hop understanding in a more robust and reliable manner.
In this work, we propose Financial Cross-Modal Multi-Hop Reasoning (FCMR), a novel benchmark designed to address the limitations of existing datasets in cross-modal multi-hop reasoning. FCMR provides multiple-choice QA samples that test the integration of facts from text, table, and chart modalities. For systematic assessment, it consists of three levels of difficulty—Easy, Medium, and Hard. As shown in Figure 1, every instance in FCMR necessitates understanding all three modalities to be answered correctly. In addition, problems at the Hard level explicitly demand cross-modal three-hop reasoning, making them more challenging (see Figure 2). Since FCMR is built using data sources from the financial domain, it is relatively free from the risk of data contamination.
Experiments on FCMR confirm that it poses challenges even for state-of-the-art MLLMs, e.g., GPT-4o and Claude 3.5 Sonnet Anthropic (2024), encouraging research efforts to develop systems capable of reasoning across multiple modalities. For analysis, we define four fine-grained procedures of cross-modal multi-hop reasoning—Planning, Modality Identification, Information Retrieval, and Information Reasoning—and probe diverse models. We reveal that models particularly struggle with the Information Retrieval phase, implying that MLLMs often fail to extract precise information from a specific modality, even when they successfully identify where the required information is located. We also present valuable findings from other analyses, such as the observation that MLLMs still have difficulty with adding negative numbers.
@@ 2.1 Cross-Modal Multi-Hop Reasoning
Meanwhile, datasets like ManyModalQA Hannan et al. (2020) and CT2C-QA Zhao et al. (2024) incorporate three modalities but lack an inherent focus on cross-modal multi-hop reasoning. MMQA Talmor et al. (2021), in contrast, deals with three modalities—text, tables, and images—and requires three-hop reasoning, setting it apart from others. It has served as the de facto standard for evaluating related methods Rajabzadeh et al. (2023); Yu et al. (2023); Luo et al. (2023); Zhang et al. (2024a); Abaskohi et al. (2024). MMCoQA Li et al. (2022) and MMCV Wang et al. (2024) have been developed as extensions of MMQA.
@@ 2.2 Limitations of MMQA
As MMQA is constructed from Wikipedia, it is vulnerable to data contamination. That is, the model being tested may already possess internalized knowledge of certain facts, reducing its dependence on the dataset’s provided input. In particular, Table 2 shows that GPT-4o can achieve reasonable performance on the most challenging part of MMQA—questions intentionally tailored for requiring a combination of information from three modalities—without relying on visual hints.Figure 9 presents an example of data leak in MMQA. This suggests that MMQA falls short of effectively measuring cross-modal multi-hop reasoning ability as it was originally intended.
@@ 3 Proposed Benchmark: FCMR
We introduce Financial Cross-Modal Multi-Hop Reasoning (FCMR), a new benchmark created to alleviate the shortcomings of MMQA and enable a more comprehensive evaluation of cross-modal multi-hop reasoning. FCMR includes three modalities—text, tables, and charts—and presents questions that entail selecting all correct statements from a set of three. The tested model must identify all true statements in the problem and is considered correct only if its final prediction is accurate.
@@ 3.1 Datset Generation Framework: CMRGen
We propose Cross-Modal Multi-Hop Reasoning Generator (CMRGen), a framework that facilitates the construction of cross-modal multi-hop reasoning datasets across various domains. CMRGen distinguishes itself from other cross-modal multi-hop dataset generation methods with its highly automated and cost-effective pipeline. Notably, while generating a single question in MMQA incurs a cost of \$0.33, our method reduces this to \$0.004 per question. Furthermore, the framework demonstrates high adaptability to various domains and offers seamless control over difficulty levels, ranging from Easy to Hard. In this study, we focus on the financial domain, where complex interactions among text, tables, and charts frequently occur, making it an ideal testbed for evaluating cross-modal multi-hop reasoning. However, the proposed pipeline is also readily applicable to domains other than finance, demonstrating its flexibility.We showcase the application of the proposed method in material science. For more details, refer to Appendix A.
@@ 3.2 Procedure of CMRGen
In the first stage, we prepare and organize data for the text, table, and chart modalities. As the origin of information, CMRGen utilizes two sources: Text Source and Table Source. For FCMR, the Text Source consists of Annual 10-K Reports collected from the SEC EDGAR database,https://www.sec.gov/search-filings/ while the Table Source is derived from Annual Simplified Financial Statements provided by WRDS Compustat.https://wrds-www.wharton.upenn.edu/pages/grid-items/compustat-global-wrds-basics/ We then filter entries that share common company entities, aligning the two sources. Finally, we construct each data instance in FCMR, comprising a document, a table, and a chart about three companies. In the next step, this instance will be supplemented with three statements serving as questions about the contents generated in this stage. Note that the chart is created by plotting specific columns from the Table Source using custom scripts, and the columns used to create the chart are removed from the table.
@@
In the final step, we apply two stages of lexical and syntactic paraphrasing using GPT-4o to enhance the diversity of expression and structure in the statements. Afterward, we conduct both LLM-based filtering (using Claude 3.5 Sonnet) and human-based filtering to ensure semantic accuracy. For Hard-level instances, human experts thoroughly review them to eliminate ambiguity and guarantee high quality. By utilizing distinct LLMs in various procedures, we aim to reduce unintended model-oriented biases.
@@ 3.3 Multiple-Choice Design
Dataset: FCMR (Hard) Image? Accuracy
Random Selection - 12.28
GPT-4o 14.71
24.37
@@ 3.4 Data Quality Control
To uphold high data quality, we implement multifaceted verification protocols.Refer to Appendix C for full details of our strategies. Specifically, we utilize Word Position Deviation (WPD) and Lexical Deviation (LD) metrics Liu and Soh (2022) to evaluate paraphrasing quality and compare these values with those from MRPC Dolan and Brockett (2005) and PAWS Zhang et al. (2019). The outcomes, presented in Table 6, confirm the superiority of our paraphrasing method. Furthermore, to verify that our dataset avoids the contamination issue identified in MMQA, we replicate the contamination experiment under the same conditions. As shown in Table 3, when the chart images are withheld, GPT-4o’s performance approximates random guessing, alleviating the risk of data contamination in FCMR. We further mitigate potential biases by equalizing modality order, statement types, and answer distributions. Figure 20 demonstrates that our benchmark is well-balanced across various perspectives. The final dataset consists of 757 Easy, 728 Medium, and 714 Hard instances, demand 100% cross-modal three-hop reasoning, setting FCMR apart from MMQA and its variants.
@@ 4.1 Experimental Setup
We evaluate a wide range of MLLMs and LLMs on FCMR under a zero-shot CoT setting, where no task-specific tuning or demonstration is provided. All models are prompted with the same template in Figure 22. Tables are represented in JSON format. For proprietary models, we employ the Claude version claude-3-5-sonnet-20241022, the GPT-4o version gpt-4o-2024-08-06., and Gemini version gemini-1.5-pro-002. We also test several open-source models: Llama variants Touvron et al. (2023), Qwen Yang and Yang (2024), MiniCPM Yao and Yu (2024), Llava Liu et al. (2024a), and ChartInstruct Masry and Shahmohammadi (2024).
@@ 4.2 Main Results
Table 4 reports the performance of various MLLM across different levels. Most open-source models perform just above random chance at the Easy level, focused on single-modal, one-hop reasoning, confirming FCMR as a challenging benchmark. Proprietary models perform better, demonstrating a remarkable gap in reasoning ability. However, at the Hard level, which necessitates full cross-modal three-hop integration, even sophisticated models, including Claude 3.5 Sonnet, achieve only around 30%. This result underscores FCMR’s challenging nature and the need for developing more advanced reasoning strategies.
@@
Experimental results indicate that open-source models with fewer than 8B parameters continue to perform comparably to random selection for tasks at the Medium and Hard levels. However, for the Easy category, they demonstrate superior performance compared to similarly sized MLLMs. Surprisingly, even advanced MLLMs such as GPT-4o and Claude 3.5 Sonnet achieve performance gains in certain cases, suggesting that their visual interpretation capabilities are still not perfect. In Section 6.2, we dive deeper into this phenomenon.
@@ 5.1 Statement-Level Analysis
Figure 5: Fine-grained stage-based analysis of three advanced MLLMs. This unique strategy discovers several intriguing findings, including that the models fails most often at (3) Information Retrieval.
@@ 5.2 Stage-Based Analysis
We manually monitor MLLMs’ inference trajectories on 40 given samples. After each fine-grained stage, we record the number of problems successfully processed by the models, forming a success history diagram in Figure 5 (b). The visualization reveals an intriguing pattern: for most samples, MLLMs fail at some stage before completing the reasoning steps for the first statement. Specifically, we observe a notable performance drop at the [Statement 1, Modality 2, (3)-(4)] stage. This suggests that while the models handle the first modality relatively well, they struggle considerably when they encounter a second modality. Interestingly, GPT-4o outperforms Claude 3.5 Sonnet in the first modality reasoning step of the first statement, but Claude surpasses GPT-4o starting from the second modality phase. As the models progress to the second statement, Claude’s performance diverges further from the others, showcasing more robust and sustained reasoning capabilities.Fine-grained model answer examples are in Figure 10.
We further explore model failures by identifying which of the four reasoning steps (1)-(4) these failures occur in, regardless of statement and modality. As illustrated in Figure 21, the most common cause of failure across MLLMs occurs at step (3) Information Retrieval—failing to extract the required information from the identified modality. The second most frequent failures arise at stage (4), Information Reasoning, where models struggle to correctly apply the retrieved information to the given conditions. Notably, Gemini 1.5 Pro exhibits a higher proportion of failures at step (4), meaning that even when it successfully retrieves information, it has difficulty reasoning over it.Examples are shown in Figure 11 and Figure 12.
@@ 6.1 Error Rate by Modality
Table 5 displays the numbers and proportions of statements Claude fails to interpret correctly, based on randomly selected 90 statements for each difficulty level.Errors unrelated to modality, such as misinterpreting conditions, are excluded from this analysis. At the Easy level, Claude frequently struggles with interpreting charts, performing noticeably worse compared to its handling of text and tables. This disparity indicates that MLLMs generally exhibit weaker capabilities in interpreting charts than in processing textual or tabular data. However, as the difficulty level increases, errors in text and tables become more prominent. This shift is likely attributed to the increased complexity of reasoning chains, which raises the likelihood of errors in interpreting text and tables.
6.2 Chart Interpretability
@@ 6.3 Error Classification and Inspection
To manually inspect Claude’s working patterns, we examine 100 error cases where it was unable to provide accurate answers. The distribution of these errors is visualized in Figure 6. Real error cases can be found in Figure 7 and Appendix D. The most common issue is again about incorrectly assessing trends in charts, such as misidentifying whether values were increasing or decreasing (35 cases). The second most frequent type of error, with 16 instances, pertains to the misidentification of top-ranked entities or overall rankings. The model also fails in 11 cases to identify entities meeting the given conditions or applies the conditions incorrectly.
A total of 17 errors stem from misinterpreting data or values, including mishandling sums, negative values, subtle differences, or column confusion. In 4 cases, the model conflates the identities of companies or entities. It also makes 8 errors in information extraction and reasoning, such as inaccurately extracting facts from text or drawing unjustified conclusions despite correct information. Lastly, 9 instances involve misunderstanding problem instructions, ignoring required modalities, or reaching illogical conclusions.

### v2

@@ 1 Introduction
Despite the recent progress in AI Touvron et al. (2023); Anthropic (2024); OpenAI (2024a), developing systems capable of human-level reasoning remains a challenge. Human cognition involves integrating information from multiple modalities to comprehend and make decisions. One domain that requires such a comprehensive understanding is finance Xie et al. (2024), where analysts must simultaneously examine textual reports, tabular data (e.g., balance sheets), and visual data (e.g., charts). For example, verifying the statement from Figure 2“The corporation, with the smallest act value in the years when the fopo value for ABBOTT LABORATORIES is below 730.5, is entitled to receive \$43 million in minimum sublease income from non-cancelable subleases.”—one must consider all the clues provided by each source, an ability we refer to as cross-modal multi-hop reasoning.
@@
Second, current benchmarks are largely focused on testing straightforward problems, such as single- and two-hop reasoning. As shown in Figure 1, MMQA Talmor et al. (2021)—one of the leading benchmarks in this field—features only about 0.8% of the queries that explicitly require three-hop cross-modal reasoning. Furthermore, in preliminary experiments, we discovered that GPT-4o OpenAI (2024a) can solve the MMQA’s most challenging problems with a 43% exact match accuracy, even without access to visual clues. This result highlights the urgent need to establish a higher standard for evaluating cross-modal multi-hop understanding in a more robust and reliable manner.
In this work, we propose Financial Cross-Modal Multi-Hop Reasoning (FCMR), a novel benchmark designed to address the limitations of existing datasets in cross-modal multi-hop reasoning. FCMR provides multiple-choice QA samples that test the integration of facts from text, tables, and charts. It consists of three levels of difficulty: Easy, Medium, and Hard. As in Figure 1, all instances in FCMR necessitate understanding three modalities to be answered correctly. In addition, problems at the Hard level explicitly demand cross-modal three-hop reasoning, making them more challenging (Figure 2). As FCMR is built using data sources from the financial domain, it is relatively free from the risk of data contamination.
Experiments on FCMR confirm that it poses challenges even for state-of-the-art MLLMs, e.g., GPT-4o and Claude 3.5 Sonnet, encouraging research efforts to develop systems capable of reasoning across multiple modalities. For analysis, we define four procedures of cross-modal multi-hop reasoning—Planning, Modality Identification, Information Retrieval, and Information Reasoning—and probe diverse models. We reveal that models particularly struggle with the Information Retrieval phase, implying that MLLMs often fail to extract information from a specific modality, even when they successfully identify where the required facts are located. We also report findings from additional analyses, including the observation that MLLMs have difficulty counting negative numbers.
@@ 2.1 Cross-Modal Multi-Hop Reasoning
Meanwhile, datasets like ManyModalQA Hannan et al. (2020), CT2C-QA Zhao et al. (2024), and MME-Finance Gan et al. (2024) incorporate three modalities but lack an inherent focus on cross-modal multi-hop reasoning.
MMQA Talmor et al. (2021), in contrast, deals with three modalities—text, tables, and images—and requires three-hop reasoning, setting it apart from others. It has served as the de facto standard for evaluating related methods Rajabzadeh et al. (2023); Yu et al. (2023a); Luo et al. (2023); Zhang et al. (2024a); Abaskohi et al. (2024). MMCoQA Li et al. (2022) and MMCV Wang et al. (2024) have been developed as extensions of MMQA.
@@ 2.2 Limitations of MMQA
As MMQA is constructed from Wikipedia, it is vulnerable to data contamination. That is, the model being tested may already possess internalized knowledge of certain facts, reducing its dependence on the dataset’s provided input. In particular, Table 2 shows that GPT-4o can achieve reasonable performance on the most challenging part of MMQA—questions intentionally tailored for requiring a combination of information from three modalities—without relying on visual hints. [^2] Figure 10 presents an example of data leak in MMQA. This suggests that MMQA falls short of effectively measuring cross-modal multi-hop reasoning ability as it was originally intended.
@@ 3 Proposed Benchmark: FCMR
We introduce Financial Cross-Modal Multi-Hop Reasoning (FCMR), a new benchmark created to alleviate the shortcomings of MMQA and enable a more comprehensive evaluation of cross-modal multi-hop reasoning. FCMR includes three modalities—text, tables, and charts—and presents questions that entail selecting all correct statements from a set of three. The tested model must identify all true statements in the problem and is considered correct only if its final prediction is accurate. This task inherently requires the model to address numerous subtasks—e.g., numerical calculation and chart interpretation—that require deep reasoning.
@@ 3.1 Datset Generation Framework: CMRGen
We propose Cross-Modal Multi-Hop Reasoning Generator (CMRGen), a framework that facilitates the construction of cross-modal multi-hop reasoning datasets across various domains. CMRGen distinguishes itself from other cross-modal multi-hop dataset generation methods with its highly automated and cost-effective pipeline. Notably, while generating a single question in MMQA incurs a cost of \$0.33, our method reduces this to \$0.004 per question. Furthermore, the framework demonstrates high adaptability to various domains and offers seamless control over difficulty levels, ranging from Easy to Hard. In this study, we focus on the financial domain, where complex interactions among text, tables, and charts frequently occur, making it an ideal testbed for evaluating cross-modal multi-hop reasoning. However, the proposed pipeline is also readily applicable to domains other than finance, demonstrating its flexibility. [^3] We showcase the application of the proposed method in material science. For more details, refer to Appendix A.
@@ 3.2 Procedure of CMRGen
In the first stage, we prepare and organize data for the text, table, and chart modalities. As the origin of information, CMRGen utilizes two sources: Text Source and Table Source. For FCMR, the Text Source consists of Annual 10-K Reports collected from the SEC EDGAR database, [^4] https://www.sec.gov/search-filings/ while the Table Source is derived from Annual Simplified Financial Statements provided by WRDS Compustat. [^5] https://wrds-www.wharton.upenn.edu/pages/grid-items/compustat-global-wrds-basics/ We then filter entries that share common company entities, aligning the two sources. Finally, we construct each data instance in FCMR, comprising a document, a table, and a chart about three companies. In the next step, this instance will be supplemented with three statements serving as questions about the contents generated in this stage. Note that the chart is created by plotting specific columns from the Table Source using custom scripts, and the columns used to create the chart are removed from the table. We create synthetic charts but ensure diversity by incorporating various chart types and styles, as well as using different libraries (see Appendix B.7).
@@
Lastly, we apply two stages of lexical and syntactic paraphrasing using GPT-4o to enhance the diversity of expressions in statements. Afterward, we conduct LLM-based filtering, using Claude 3.5 Sonnet, to ensure semantic accuracy. By employing separate LLMs in each step, we aim to mitigate model-oriented biases. For Hard-level instances, human experts optionally review them to further refine quality.
@@ 3.4 Data Quality Control
To uphold high data quality, we implement multifaceted verification protocols. [^6] Refer to Appendix C for full details of our strategies. Specifically, we utilize Word Position Deviation (WPD) and Lexical Deviation (LD) metrics Liu and Soh (2022) to evaluate paraphrasing quality and compare these values with those from MRPC Dolan and Brockett (2005) and PAWS Zhang et al. (2019). The outcomes, presented in Table 6, confirm the superiority of our paraphrasing method. Furthermore, to verify that our dataset avoids the contamination issue identified in MMQA, we replicate the contamination experiment under the same conditions. As shown in Table 8, when the chart images are withheld, GPT-4o’s performance approximates random guessing, alleviating the risk of data contamination in FCMR. We further mitigate potential biases by equalizing modality order, statement types, and answer distributions. Figure 21 demonstrates that our benchmark is well-balanced across various perspectives. The final dataset consists of 757 Easy, 728 Medium, and 714 Hard instances, demand 100% cross-modal three-hop reasoning, setting FCMR apart from MMQA and its variants.
@@ 4.1 Experimental Setup
We evaluate a wide range of MLLMs and LLMs on FCMR under a zero-shot CoT setting, where no task-specific tuning or demonstration is provided. This setting reflects common use cases for MLLMs, ensuring an unbiased model evaluation and capturing overall performance trends. All models are prompted with the same template in Figure 23. Tables are represented in JSON format. For proprietary models, we employ the Claude version claude-3-5-sonnet-20241022, the GPT-4o version gpt-4o-2024-08-06., and Gemini version gemini-1.5-pro-002. We also test several open-source models: Llama variants Touvron et al. (2023), Qwen Yang and Yang (2024), MiniCPM Yao and Yu (2024), Llava Liu et al. (2024a), and ChartInstruct Masry and Shahmohammadi (2024).
@@ 4.2 Main Results
Table 3 reports the performance of various MLLM across different levels. Most open-source models perform just above random chance at the Easy level, focused on single-modal, one-hop reasoning, confirming FCMR as a challenging benchmark. Proprietary models perform better, demonstrating a remarkable gap in reasoning ability. However, at the Hard level, which necessitates full cross-modal three-hop integration, even sophisticated models, including Claude 3.5 Sonnet, achieve only around 30%. [^7] Recently, deep reasoning models, e.g., OpenAI’s o1 OpenAI (2024b), have emerged. While they are not directly related to multimodal functionality, we also evaluate them at the Hard level. As a result, we confirm that o1 achieves 43% accuracy, indicating they are still far from perfect (refer to Table 9). This result underscores FCMR’s challenging nature and the need for developing more advanced reasoning strategies.
@@
Experimental results indicate that open-source models with fewer than 8B parameters continue to perform comparably to random selection for tasks at the Medium and Hard levels. However, for the Easy category, they demonstrate superior performance compared to similarly sized MLLMs. Surprisingly, even advanced MLLMs such as GPT-4o and Claude 3.5 Sonnet achieve performance gains in certain cases, suggesting that their visual interpretation capabilities are still not perfect. In Section E, we dive deeper into this phenomenon.
@@ 5.1 Statement-Level Analysis
Figure 5: Fine-grained stage-based analysis of three advanced MLLMs. This unique strategy discovers several intriguing findings, including that the models fail most often at (3) Information Retrieval.
@@ 5.2 Stage-Based Analysis
We manually monitor MLLMs’ inference trajectories on 40 given samples. After each fine-grained stage, we record the number of problems successfully processed by the models, forming a success history diagram in Figure 5 (b). The visualization reveals an intriguing pattern: for most samples, MLLMs fail at some stage before completing the reasoning steps for the first statement. Specifically, we observe a notable performance drop at the [Statement 1, Modality 2, (3)-(4)] stage. This suggests that while the models handle the first modality relatively well, they struggle considerably when they encounter a second modality. Interestingly, GPT-4o outperforms Claude 3.5 Sonnet in the first modality reasoning step of the first statement, but Claude surpasses GPT-4o starting from the second modality phase. As the models progress to the second statement, Claude’s performance diverges further from the others, showcasing more robust and sustained reasoning capabilities. [^8] Fine-grained model answer examples are in Figure 11.
We further explore model failures by identifying which of the four reasoning steps (1)-(4) these failures occur in, regardless of statement and modality. As illustrated in Figure 22, the most common cause of failure across MLLMs occurs at step (3) Information Retrieval—failing to extract the required information from the identified modality. The second most frequent failures arise at stage (4), Information Reasoning, where models struggle to correctly apply the retrieved information to the given conditions. Notably, Gemini 1.5 Pro exhibits a higher proportion of failures at step (4), meaning that even when it successfully retrieves information, it has difficulty reasoning over it. [^9] Examples are shown in Figure 12 and Figure 13.
@@ 6.1 Error Rate by Modality
Table 4 displays the numbers and proportions of statements Claude fails to interpret correctly, based on random 90 statements for each difficulty level. [^10] Errors unrelated to modality, such as misinterpreting conditions, are excluded from this analysis. At the Easy level, Claude often struggles with analyzing charts, performing noticeably worse compared to its handling of text and tables. This disparity indicates that MLLMs exhibit weaker capabilities in interpreting charts than in processing textual or tabular data (see Appendix E for details). However, as the difficulty level increases, errors in text and tables become more prominent. This shift is likely attributed to the increased complexity of reasoning chains, which raises the likelihood of errors in processing text and tables.
@@ 6.2 Error Classification and Inspection
To manually inspect Claude’s working patterns, we examine 100 error cases where it was unable to provide accurate answers. The distribution of these errors is visualized in Figure 6. Real error cases can be found in Figure 7 and Appendix D.
The most common issue involves misinterpreting trends in charts (35 cases). The second most frequent error type (16 samples) pertains to the misidentification of top-ranked entities or overall rankings. The model fails in 11 cases to correctly identify entities that meet the given conditions or misapplies the conditions. 17 errors arise from misinterpreting data, including sums, negative values, subtle differences, and column confusion. 4 involve conflating company or entity identities. The model also makes 8 errors in information extraction and reasoning, such as misreading facts or drawing unjustified conclusions. Lastly, 9 cases stem from misinterpreting instructions, ignoring required modalities, or making illogical inferences.
@@
6.3 Preliminary Task Optimization
To explore ways to enhance MLLM performance on FCMR, we conduct preliminary experiments in modality integration and prompt optimization. Table 5 presents the results of applying three techniques—Modality Integration, 4-Stage Reasoning, and Self-Refine—to 100 Hard samples.111111 Modality Integration: convert all forms of knowledge into text representations, employing chart captioning and table linearization Yu et al. (2023b); Luo et al. (2023); Zhi Lim et al. (2024). 4-Stage Reasoning: derives insights from Section 5.2 and explicitly mentions four reasoning stages in the prompt. Self-Refine: instructs the model to iteratively refine its own answer Madaan et al. (2023). The prompts for each method are shown in Figure 25, Figure 26, and Figure 27, respectively. The results confirm each method’s contribution to improved accuracy, verifying their effectiveness. This method also mitigates failure rates in both the Information Retrieval and Reasoning stages mentioned in Section 5.2 (see Figure 9). However, considering that the combination of all three methods saturates at 46% accuracy, future work is encouraged to develop dedicated approaches for FCMR.
Method (Claude 3.5 Sonnet; tested on 100 Hard cases) Acc. (%)
Zero-Shot COT (Baseline) 32
Modality Integration 39
Modality Integration + 4-Stage Reasoning 42
Modality Integration + 4-Stage Reasoning + Self-Refine 46
Table 5: Preliminary task optimization results.